OpenStudy (anonymous):

what is 3 1/2 - 1 1/5?

OpenStudy (anonymous):

I don't know what it is!

OpenStudy (anonymous):

3 1/2 - 1 1/5 can be solved as.. 5/2 - 6/5 = 25/10 - 12/10 = 13/10

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

u're welcome

OpenStudy (anonymous):

3 1/2 - 1 1/5 = (3 + 1/2) - (1 + 1/5) =\[\frac{7}{2}+\frac{6}{5}=\frac{47}{10} \]